# WisdoM[Uncaptioned image]: Improving Multimodal Sentiment Analysis by Fusing Contextual World Knowledge

### 12 Jan 2024 v1

@@ Abstract
Sentiment analysis is rapidly advancing by utilizing various data modalities (e.g., text, image). However, most previous works relied on superficial information, neglecting the incorporation of contextual world knowledge (e.g., background information derived from but beyond the given image and text pairs) and thereby restricting their ability to achieve better multimodal sentiment analysis. In this paper, we proposed a plug-in framework named WisdoM[Uncaptioned image] , designed to leverage contextual world knowledge induced from the large vision-language models (LVLMs) for enhanced multimodal sentiment analysis. WisdoM[Uncaptioned image] utilizes a LVLM to comprehensively analyze both images and corresponding sentences, simultaneously generating pertinent context. To reduce the noise in the context, we also introduce a training-free Contextual Fusion mechanism. Experimental results across diverse granularities of multimodal sentiment analysis tasks consistently demonstrate that our approach has substantial improvements (brings an average +1.89 F1 score among five advanced methods) over several state-of-the-art methods. Code will be released.
@@ 1 Introduction
Sentiment analysis (SA, Medhat et al., 2014; Wankhade et al., 2022) focuses on identifying human sentiment polarity Das and Singh (2023). With the development of the Internet, the way people express their sentiments is not limited to only text, but also includes multimodal data (e.g., images). It is hard to detect the sentiment due to the short and informal nature of the sentence sometimes, but the associated image may help reflect people’s sentiment. How to use the information between different modalities becomes the main problem of multimodal sentiment analysis (MSA).
@@
To address this problem, the relevant modal fusion strategies are carefully designed for MSA tasks of different granularity. For aspect-level MSA tasks, existing methods estimate whether the image contains information conducive to multimodal sentiment analysis by learning the relationships between image and text Ju et al. (2021); Ling et al. (2022); Yang et al. (2022); Zhou et al. (2023). For sentence-level MSA tasks, existing methods primary focus lies in the fusion at the feature level and decision level Hazarika et al. (2020); Yu et al. (2021); Han et al. (2021); Guo et al. (2022)
@@
Figure 1: The simple schematic of our method. The sentiment polarity of Aleppo is negative, which is hard to directly predict by existing methods while our WisdoMRefer to caption predict correctly via incorporating context generated by the world knowledge-rich LVLMs.
Despite empirically success of above fusion techniques, there lacks investigations about improving MSA with (easy-to-acquired) world knowledge. This raises the following question: Could world knowledge boost MSA? For instance, in Fig. 1, given a comparative image at different periods alongside a sentence. Question: What’s the sentiment polarity of “Aleppo”? Current state-of-the-art (SOTA) multimodal sentiment analysis model (MSAM) Zhou et al. (2023) can only discern that the image is a comparative of Aleppo at various period. The predicted sentiment polarity is neutral. This is because the model lacks the knowledge of the Aleppo. In fact, human would associate Aleppo with being a city in Syria, and in conjunction with the image, we might infer connections to the Syrian war. Consequently, the deduction would be that the intent is to express the severe devastation experienced by the city of Aleppo due to the Syrian war, highlighting a stark contrast between its once-prosperous city and its current dilapidation. Hence, the sentiment towards Aleppo is negative.
@@
Inspired by the human knowledge-guided in learning, we propose a plug-in framework WisdoM[Uncaptioned image] , utilizing contextual world knowledge from large vision-language models for MSA. Specifically, the WisdoM[Uncaptioned image] contains three stages: ❶Prompt Templates Generation, ❷Context Generation, and ❸Contextual Fusion. In stage 1, we ask ChatGPT [^1] https://chat.openai.com to generate prompt templates which are used to construct instructions for stage 2. The goal in stage 2 is to enables the large vision-language model (LVLM) Liu et al. (2023b) to generate contextual information based on the provided image and sentence. Note that, we refer to this contextual information as context. However, the context were found to be noisy in MSA. We further introduce a training-free Contextual Fusion that incorporating context for hard sample in stage 3.
@@
We validated our WisdoM[Uncaptioned image] on several benchmarks including Twitter2015, Twitter2017 Yu and Jiang (2019) and MSED Jia et al. (2022) over several models: LLaVA-v1.5 Liu et al. (2023a), MMICL Zhao et al. (2023), mPLUG-Owl2 Ye et al. (2023), AoM Zhou et al. (2023), ALMT Zhang et al. (2023). Experiments demonstrate the effectiveness and universality of our approach. Extensive analysis confirms that WisdoM[Uncaptioned image] effectively enhances the ability of MSAM on MSA tasks.
Contributions
Our main contributions are:
@@
  • We propose a plug-in framework WisdoM[Uncaptioned image] , leveraging the LVLM to generate explicit contextual world knowledge, to wisely enhance the multimodal sentiment analysis ability.
@@
  • We conduct experiments on three multimodal sentiment analysis benchmarks upon several advanced VLP models, showing that WisdoM[Uncaptioned image] could bring consistent and significant improvements (up to +5.8 F1 score).
@@ 2.1 Multimodal Sentiment Analysis
Different from traditional text-based sentiment analysis task Hussein (2018), multimodal sentiment analysis (MSA) task Soleymani et al. (2017) integrates more modalities simultaneously, e.g., image and speech, to obtain accurate sentiment labels. Numerous advanced models have been proposed, covering different levels of granularity, such as sentence and aspect Jia et al. (2022). The core of most MSA studies lies in developing better fusion algorithms across modalities.
Sentence-level MSA
Wang et al. (2014) introduce a cross-media bag-of-words approach in the joint image-text framework for microblog images. You et al. (2015) employ deep neural networks to conduct a combined textual and visual sentiment analysis. Zhao et al. (2019) present a sentiment analysis model driven by image-text consistency, exploring correlations between image and text data in a movie review dataset. Li et al. (2020) propose a ConvTransformer, a multi-window convolutional transformer model that combines Transformer Vaswani et al. (2017) and CNN to identify joint sentiment polarity. Das and Singh (2022) propose a multi-stage multimodal sentiment analysis model for the Assamese language, utilizing a deep learning architecture that concurrently exploits textual and visual cues to determine sentiment. Zhang et al. (2023) propose an Adaptive Language-guided Multimodal Transformer (ALMT), which incorporates an Adaptive Hyper-modality Learning (AHL) module to learn an irrelevance/conflict-suppressing representation from different modalities under the guidance of language features at different scales.
Aspect-level MSA
Multimodal aspect-oriented sentiment classification combines both multimodal sentiment analysis Bezemer and Jewitt (2010) and aspect-oriented sentiment classification Wang et al. (2022); Zhong et al. (2023a), typically applied to multimedia content sourced from online social media Khan and Fu (2021). Yu and Jiang (2019) introduce TomBERT, leveraging two multimodal tweet datasets with target annotations, and proposed an aspect-oriented multimodal BERT model. TomBERT extends the baseline BERT architecture Kenton and Toutanova (2019) by incorporating target-sensitive visual attention and augmenting self-attention layers to capture inter-modality dynamics. Khan and Fu (2021) propose a two-stream model employing an object-aware transformer for image translation in the input space, followed by a single-pass non-autoregressive generation approach Wu et al. (2020); Ding et al. (2021).
Although existing approaches relying on sophisticated fusion techniques have achieved remarkable performance in MSA, their limitation lies in relying superficial information for fusion, without incorporating contextual world knowledge.
@@ 2.2 Large Vision-Language Models
Large Vision-Language Models (LVLMs) are becoming a fundamental tool for general task solver Li et al. (2023); Fu et al. (2023); Liu et al. (2023b); Zhao et al. (2023); Ye et al. (2023); Dai et al. (2023); Zhu et al. (2023); Alayrac et al. (2022); Chen et al. (2023). Liu et al. (2023a) introduce LLaVA, which establishes a connection between the pre-trained CLIP ViT-L/14 visual encoder Dosovitskiy et al. (2020) and the large language model Vicuna Chiang et al. (2023) through a straightforward projection matrix. The model is trained using a two-stage instruction-tuning procedure. Addressing the challenge of complex multi-modal prompts with multiple images, Zhao et al. (2023) propose MMICL, considering both model and data perspectives to enhance the understanding of LVLMs. Ye et al. (2023) propose mPUG-Owl2, a versatile multi-modal large language model that effectively utilizes modality collaboration to enhance performance in both text and multi-modal tasks. mPLUG-Owl2 employs a modularized network design, with the language decoder serving as a universal interface for managing various modalities.
Here, we explore the utility of existing LVLMs in generating contextual world knowledge for imporving multimodal sentiment analysis. Additionally, we introduce a training-free module, Contextual Fusion, to reduce the noise in the context.
@@
Figure 3: Detailed illustration of our proposed schema WisdoMRefer to caption with a running example. ❶ Using ChatGPT to provide prompt templates. ❷ We then prompt LVLMs to generate context using the prompt templates with image and sentence. ❸ A training-free mechanism Contextual Fusion mitigate the noise in the context.
@@ 3 Preliminary
In this section, we first describe the notation of the MSA. We then review two typical frameworks to modeling the MSA tasks, where we experiment our schema upon them: task-specific framework Zhou et al. (2023); Ling et al. (2022); Ju et al. (2021) and general-purpose framework Li et al. (2023).
Notation
@@
Task-Specific Framework
@@
General-Purpose Framework
To verify that our WisdoM works well on arbitrary architectures, we also apply WisdoM to LVLMs. We follow Wang et al. (2023) to construct the task instructions $I_{aspect}$ and $I_{sentence}$ for each task to elicit its ability to the corresponding task. The task instructions are presented as single-choice questions with well-formatted options (Shown in Fig. 2). The LVLMs can be seen as a sentiment classifier $f(\cdot)$. Therefore, the aspect-level task and sentence-level task can be formulated as $(y_{i},P_{i})=f(I_{aspect},v_{i},s_{i},a_{i})$ and $(y_{i},P_{i})=f(I_{sentence},v_{i},s_{i})$ respectively.
@@ 4 Methodology
Overview
Fig. 3 illustrates the overview of our method following three stages. In the ❶Prompt Templates Generation, we use ChatGPT to provide prompt templates. These prompt templates are fed into the LVLM with sentence $s_{i}$ and image $v_{i}$ to generate context, also called the ❷Context Generation. During ❸Contextual Fusion, we first compute the confidence, determining if the sample is uncertain (referred to as a hard sample). For hard samples, we fuse the predicted probability $P_{i}$ with $\hat{P}_{i}$ which obtained by incorporating context. Otherwise, we use $P_{i}$ as the final prediction.
@@ 4.1 Stage 1: Prompt Templates Generation
The main purpose of this stage is to design the prompt templates used to generate the context, so that the LVLM can better understand our intention and thus provide a more comprehensive contextual world knowledge. Inspired by Jiao et al. (2023); Zhong et al. (2023b), we ask ChatGPT to provide the appropriate prompt templates. The prompt templates provided by ChatGPT consider world knowledge of different perspectives, including history, social and cutural, etc. We insert a “Sentence: [x]” at the end of the prompt template to place the input sentence $s_{i}$. The example of prompt templates are shown in Appendix A.2.
4.2 Stage 2: Context Generation
In the context generation stage, prompt templates in ❶ are used to generate context which explicit world knowledge based on given the image $v_{i}$ and sentence $s_{i}$ by LVLMs Liu et al. (2023b); Ye et al. (2023). Specifically, we construct instruction by replacing the “[x]” in the prompt template with the sentence $s_{i}$. In addition, different LVLMs require a special token to indicate where the image $v_{i}$ is inserted. Taking LLaVA Liu et al. (2023b) as an example, we insert a special token “<image>” at the beginning of the instruction.
4.3 Stage 3: Contextual Fusion
We use the sentence-level task as an example. After obtained the context, we can intuitively use predicted sentiment polarity $\hat{y}_{i}$ obtained by incorporating context, i.e., $(\hat{y}_{i},\hat{P}_{i})=f(v_{i},s_{i},context)$. However, the context may contains irrelevant information that could disturb performance. Therefore, we first determine hard sample and then fuse the $P_{i}$ and $\hat{P}_{i}$ in hard sample.
Determine Hard Sample
Inspired by Zhang et al. (2021), we found that the ambiguous hard sample are commonly found around boundary areas of the sentiment polarity (e.g., the boundary areas of negative and neutral). Therefore, given a sample $m_{i}$, we only consider the difference $\delta_{i}$ between the highest and the second highest probabilities to determine whether is a hard sample:
$\delta_{i}=2\times max(P_{i})+min(P_{i})-1.$ (1)
Fusion with Context
$\tilde{P}_{i}=P_{i}+\beta\cdot(\hat{P}_{i}-P_{i}),$ (2)
where $\beta$ is an interpolation coefficient. Intuitively, $(\hat{P}_{i}-P_{i})$ represent the information incorporating by context. $\beta$ is used to control the proportion of information introduced in context. When $\beta\rightarrow 0$, the effect brought by context is completely ignored and vice versa. Note that, we use $(y_{i},P_{i})$ as the final prediction when $m_{i}\notin\mathcal{V}_{hard}$.
5 Experiment
@@
TomBERT Yu and Jiang (2019) 77.2 71.8 70.5 68.0
CapTrBERT Khan and Fu (2021) 78.0 73.2 72.3 70.2

### 20 Feb 2024 v2

@@ Abstract
Sentiment analysis is rapidly advancing by utilizing various data modalities (e.g., text, image). However, most previous works relied on superficial information, neglecting the incorporation of contextual world knowledge (e.g., background information derived from but beyond the given image and text pairs) and thereby restricting their ability to achieve better multimodal sentiment analysis (MSA). In this paper, we proposed a plug-in framework named WisdoM[Uncaptioned image] , to leverage the contextual world knowledge induced from the large vision-language models (LVLMs) for enhanced MSA. WisdoM utilizes LVLMs to comprehensively analyze both images and corresponding texts, simultaneously generating pertinent context. To reduce the noise in the context, we also introduce a training-free contextual fusion mechanism. Experiments across diverse granularities of MSA tasks consistently demonstrate that our approach has substantial improvements (brings an average +1.96% F1 score among five advanced methods) over several state-of-the-art methods. The code will be released.
[…]
shadows
@@ 1 Introduction
Sentiment analysis (SA, Medhat et al., 2014; Wankhade et al., 2022) task focuses on identifying human sentiment polarity Das and Singh (2023). With the development of the Internet, the way people express their sentiments is not limited to text, but also includes multimodal data (e.g., images). Detecting the sentiments of such information is challenging because of their short, informal nature, but utilizing the paired images can offer valuable insights. Therefore, how to make accurate sentiment classification by effectively marring modalities is at the core of multimodal sentiment analysis (MSA, Majumder et al., 2018; Wang et al., 2020).
@@
Recent studies that improve MSA by carefully designing fusion strategies can be categorized into two types: 1) for aspect-level MSA, existing methods Ju et al. (2021); Ling et al. (2022); Yang et al. (2022); Zhou et al. (2023) estimate whether the image contains conducive information by learning the relationship between images and text, and locally fuse the aspect-aware images; 2) for sentence-level MSA, researchers primarily focus on the global fusion at feature-level and decision-level Hazarika et al. (2020); Yu et al. (2021); Han et al. (2021); Guo et al. (2022).
@@
Figure 1: The simple schematic of our method. The sentiment polarity of Aleppo is negative, which is hard to directly predict by existing methods while our WisdoMRefer to caption predicts correctly via incorporating context generated by the world knowledge-rich LVLMs.
Despite their empirical success, the above studies only consider the superficial information111only reflects the surface or literal information without considering their deep (e.g., historical and cultural) meaning. between image and text (See the case “Aleppo” in Fig. 1), and sometimes it is difficult to predict the true polarity without their background world knowledge (“Aleppo has been severely affected by the ongoing Syrian Civil War…” induced from the large vision-language models.). This raises the following question: Could world knowledge boost MSA? Take the test case in Fig. 1 as an instance, given a comparative image at different periods alongside a sentence, it is required to answer the question: What’s the sentiment polarity of “Aleppo”?. We employ the current state-of-the-art (SOTA) MSA model Zhou et al. (2023) as the backbone. As expected, even the SOTA model gives the wrong prediction (i.e., “neutral” rather than the groundtruth– “negative”) because of lacking the deeper knowledge of “Aleppo” (which is a city in Syria, in conjunction with the image, we might infer that the difference between the before and after of this city is caused by the Syrian war). Therefore, employing world knowledge is essential.
@@
Correspondingly, we propose a plug-and-play framework to utilize the contextual world knowledge (simply induced from the large vision-language models) to complement the existing text-image pair with only superficial information, namely WisdoM[Uncaptioned image] . In particular, our WisdoM follows a three stages process: 1 Prompt Templates Generation, 2 Context Generation, and 3 Contextual Fusion. In stage 1, we ask language models (e.g., ChatGPT [^2] https://chat.openai.com) to generate prompt templates (See the Prompt Template “Give you a sentence and image, you can …” in Fig. 1) which are used to construct instructions for stage 2. Then, we employ the advanced large vision-language model (LVLM, e.g., LLaVA Liu et al., 2023b in stage 2 to generate the contextual information (See the Context “Aleppo, … ongoing Syrian Civil War…” in Fig. 1) based on the provided image and sentence. Note that, we refer to this contextual information as context. Due to the noisy nature of the derived context, we further introduce a training-free Contextual Fusion mechanism, in stage 3, to wisely incorporate the context for hard samples.
@@
We validated our WisdoM on several benchmarks including Twitter2015, Twitter2017 Yu and Jiang (2019) and MSED Jia et al. (2022) over several models: LLaVA-v1.5 Liu et al. (2023a), MMICL Zhao et al. (2023), mPLUG-Owl2 Ye et al. (2023), AoM Zhou et al. (2023), ALMT Zhang et al. (2023b). Experiments demonstrate the effectiveness and universality of our approach, and extensive analyses provide insights into when and how our method works. Our main contributions are:
@@
  • We propose a plug-in framework WisdoM, leveraging the LVLM to generate explicit contextual world knowledge, to enhance the multimodal sentiment analysis ability.
@@
  • Experiments on three MSA benchmarks upon several advanced LVLMs, show that WisdoM brings consistent and significant improvements (up to +6.3% F1 score).
@@ 2.1 Multimodal Sentiment Analysis
Multimodal Sentiment Analysis (MSA), diverging from conventional text-based approaches Hussein (2018), incorporates diverse modalities (e.g., image, speech) to enhance sentiment classification accuracy Soleymani et al. (2017). Numerous advanced models have been proposed, covering different levels of granularity, such as sentence and aspect:
Sentence-Level MSA.
Wang et al. (2014) integrate images and text for microblog analysis. You et al. (2015) employ deep neural networks for textual and visual sentiment analysis. Zhao et al. (2019) explore image-text correlations in movie reviews. Li et al. (2020) propose a ConvTransformer, blending Transformer Vaswani et al. (2017) and CNN technologies for sentiment analysis. Das and Singh (2022) propose a multi-stage multimodal method for the Assamese language, leveraging both text and images. Zhang et al. (2023b) present an advanced model, ALMT, enhancing multimodal analysis with a focus on language-guided features to handle irrelevant or conflicting data across different modalities.
Aspect-Level MSA.
Yu and Jiang (2019) introduce TomBERT, leveraging two multimodal tweet datasets with target annotations, and proposed an aspect-oriented multimodal BERT model. Khan and Fu (2021) propose a two-stream model employing an object-aware transformer for image translation in the input space, followed by a single-pass non-autoregressive generation approach Wu et al. (2020); Ding et al. (2021). Zhou et al. (2023) introduce an aspect-oriented network, AoM, designed to reduce visual and textual distractions arising from intricate image-text interactions.
Although existing approaches relying on sophisticated fusion techniques have achieved remarkable performance in MSA, their limitation lies in relying on superficial information for fusion, without incorporating contextual world knowledge.
@@ 2.2 Large Vision-Language Models
Large Vision-Language Models (LVLMs) are becoming a fundamental tool for solving general tasks Li et al. (2023); Fu et al. (2023); Liu et al. (2023b); Zhao et al. (2023); Ye et al. (2023); Dai et al. (2023); Zhu et al. (2023); Alayrac et al. (2022); Chen et al. (2023). Liu et al. (2023a) introduce LLaVA, which connects the CLIP ViT-L/14 visual encoder Dosovitskiy et al. (2020) with the large language model Vicuna Chiang et al. (2023b) through a simple projection matrix, utilizing a two-stage instruction-tuning method. Zhao et al. (2023) propose MMICL to tackle the complexity of multi-modal prompts, focusing on improving LVLMs from both model and data viewpoints. Ye et al. (2023) propose mPUG-Owl2, a multi-modal language model designed for enhanced collaboration between modalities. It features a modular network where the language decoder acts as a universal interface for different modalities.
Here, we leverage LVLMs to enhance multimodal sentiment analysis by generating relevant world knowledge. Additionally, we introduce a new training-free module named Contextual Fusion, designed to minimize noise in the context.
2.3 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) enhances Language Models by incorporating retrieved text, significantly improving performance in knowledge-based tasks, applicable in both fine-tuned and off-the-shelf scenarios Gao et al. (2023); Gupta et al. (2024). Traditional RAG Lewis et al. (2020), also known as Naive RAG, incorporates retrieval content to aid generation but faces key challenges: 1) varying retrieval quality, 2) generation of responses prone to inaccuracies, and 3) difficulties in coherently integrating retrieved-context with current tasks. To overcome the limitations of Naive RAG, advanced methods introduce more contextually rich information during inference. The DSP framework Khattab et al. (2022) facilitates an intricate exchange between frozen LMs and retrieval models, improving context richness, while PKG Luo et al. (2023) allows LLMs to retrieve relevant information for complex tasks without altering their parameters. The working mechanism of WisdoM is similar to RAG, but different at the following aspects: ① WisdoM utilizes LVLM to generate world knowledge to provide coherent and accurate context rather than retrieval, ② WisdoM incorporates a contextual fusion mechanism to diminish noise within the context. For additional experimental analysis and discussion, please refer to § 5.3.
@@ 2.3 Retrieval-Augmented Generation
Figure 3: Detailed illustration of our proposed schema WisdoMRefer to caption with a running example. 1 Using ChatGPT to provide prompt templates. 2 We then prompt LVLMs to generate context using the prompt templates with image and sentence. 3 A training-free mechanism Contextual Fusion mitigates the noise in the context.
@@ 3 Preliminary
We first describe the notation of the MSA, then review two typical frameworks for modelling the MSA tasks, where we experiment with our schema upon them: task-specific framework Zhou et al. (2023); Zhang et al. (2023b) and general-purpose framework Liu et al. (2023a); Ye et al. (2023).
Notation.
@@
Task-Specific Framework.
@@
General-Purpose Framework.
To verify that our WisdoM works well on arbitrary architectures, we also apply WisdoM to general-purpose LVLMs. We follow Wang et al. (2023b) to construct the task instructions $I_{aspect}$ and $I_{sentence}$ for each task to elicit its ability to the corresponding task. The task instructions are presented as single-choice questions with well-formatted options (Shown in Fig. 2). The LVLMs can be seen as a sentiment classifier $f(\cdot)$. Therefore, the aspect-level task and sentence-level task can be formulated as $(y_{i},P_{i})=f(I_{aspect},v_{i},s_{i},a_{i})$ and $(y_{i},P_{i})=f(I_{sentence},v_{i},s_{i})$ respectively.
@@ 4 Methodology
Overview.
Fig. 3 illustrates the overview of our method following three stages. In the 1 Prompt Templates Generation, we use large language models, particularly ChatGPT, to provide prompt templates. These prompt templates are fed into the LVLM with sentence $s_{i}$ and image $v_{i}$ to generate context, also called the 2 Context Generation. During 3 Contextual Fusion, we first compute the confidence, determining if the sample is uncertain (referred to as a hard sample). For hard samples, we fuse the predicted probability $P_{i}$ with $\hat{P}_{i}$ which is obtained by incorporating context. Otherwise, we use $P_{i}$ as the final prediction.
@@ 4.1 Stage 1: Prompt Templates Generation
The main purpose of this stage is to design the prompt templates used to generate the context, so that the LVLM can better understand our intention and thus provide a more comprehensive contextual world knowledge. Inspired by Jiao et al. (2023); Zhong et al. (2023), we ask ChatGPT to provide the appropriate prompt templates. The prompt templates provided by ChatGPT consider world knowledge of different perspectives, including history, social and cultural, etc. We insert a “Sentence: [x]” at the end of the prompt template to place the input sentence $s_{i}$. The example of prompt templates are shown in Appendix A.5.
@@
TomBERT Yu and Jiang (2019) 77.2 71.8 70.3 68.0
CapTrBERT Khan and Fu (2021) 77.9 73.9 72.3 70.2
@@
4.2 Stage 2: Context Generation
In the context generation stage, prompt templates in 1 are used to generate context that explicitly incorporates world knowledge based on given the image $v_{i}$ and sentence $s_{i}$ by LVLMs Liu et al. (2023b); Ye et al. (2023). Specifically, we construct instruction by replacing the “[x]” in the prompt template with the sentence $s_{i}$. In addition, different LVLMs require a special token to indicate where the image $v_{i}$ is inserted. Taking LLaVA Liu et al. (2023b) as an example, we insert a special token “<image>” at the beginning of the instruction.
4.3 Stage 3: Contextual Fusion
We use the sentence-level task as an example. After obtaining the context, we can intuitively use predicted sentiment polarity $\hat{y}_{i}$ obtained by incorporating context, i.e., $(\hat{y}_{i},\hat{P}_{i})=f(v_{i},s_{i},context)$. However, the context may contain irrelevant information that could disturb performance. Therefore, we first determine the hard samples and then fuse the $P_{i}$ and $\hat{P}_{i}$ in the hard sample.
Determining the Hard Samples.
Inspired by Zhang et al. (2021), we found that the ambiguous hard sample is commonly found around boundary areas of the sentiment polarity (e.g., the boundary areas of negative and neutral). Therefore, given a sample $m_{i}$, we only consider the difference $\delta_{i}$ between the highest and the second highest probabilities to determine whether it is a hard sample:
$\delta_{i}=2\times max(P_{i})+min(P_{i})-1.$ (1)
Fusion with Context.
$\tilde{P}_{i}=P_{i}+\beta\cdot(\hat{P}_{i}-P_{i}),$ (2)
where $\beta$ is an interpolation coefficient. Intuitively, $(\hat{P}_{i}-P_{i})$ represents the information incorporating by context. $\beta$ is used to control the proportion of information introduced in context. When $\beta\rightarrow 0$, the effect brought by context is completely ignored and vice versa. Note that, we use $(y_{i},P_{i})$ as the final prediction when $m_{i}\notin\mathcal{V}_{hard}$. We study the impact of $\alpha$ and $\beta$ in Appendix B.1.
5 Experiments